OpenStudy (anonymous):

onsider the line that passes through the point (3, -2) and has a slope of 2. Part 1: Write the equation of this line using point-slope form. (2 points) Part 2: Using your equation from part 1, rewrite this equation in slope-intercept form. Make sure to show all of your work. (2 points) Part 3: Using your equation from part 2, rewrite this equation in standard form. Make sure to show all of your work. (2 points)

OpenStudy (anonymous):

@AbdullahM Last thing im going to need help with :3

OpenStudy (abdullahm):

Point-Slope form: \(\sf y – y_1 = m(x – x_1)\)

OpenStudy (abdullahm):

Can you do part 1? (:

OpenStudy (abdullahm):

Slope-intercept form: \(\sf y=mx+b\)

OpenStudy (anonymous):

how would that work? we onl have one point

OpenStudy (abdullahm):

\(\sf (3,-2)\) is in the form of \(\sf (x_1,y_1)\)

OpenStudy (abdullahm):

Just subsitute the values (: